Saints Company sells 50,000 units at $25 per unit. Fixed costs are $70,000, and income from operations is $380,000. Determine: A) Variable Cost Per Unit, B) Unit Contribution Margin, C) Contribution Margin Ratio.
Answer #1

Contribution Margin=Fixed cost+Income from operations

=(70,000+380,000)=$450,000

Hence Contribution Margin per unit=(450,00/50,000)=$9 per unit

Contribution Margin ratio=Contribution Margin/Sales

=9/25=36%

Contribution margin=Sales-Variable cost

Variable cost=(25-9)=$16 per unit

Hence:

A) Variable Cost Per Unit=$16 per unit

B) Unit Contribution Margin=$9 per unit

C) Contribution Margin Ratio=36%